### v1

# A unified framework for coarse grained molecular dynamics of proteins

Abstract
Understanding protein dynamics is crucial for elucidating their biological functions. While all-atom molecular dynamics (MD) simulations provide detailed information, coarse-grained (CG) MD simulations capture the essential collective motions of proteins at significantly lower computational cost. In this article, we present a unified framework for coarse-grained molecular dynamics simulation of proteins. Our approach utilizes a tree-structured representation of collective variables, enabling reconstruction of protein Cartesian coordinates with high fidelity. The force field is constructed using a deep neural network trained on trajectories generated from conventional all-atom MD simulations. We demonstrate the framework’s effectiveness using the 168-amino protein target T1027 from CASP14. Statistical distributions of the collective variables and time series of root mean square deviation (RMSD) obtained from our coarse-grained simulations closely resemble those from all-atom MD simulations. This method is not only useful for studying the movements of complex proteins, but also has the potential to be adapted for simulating other biomolecules like DNA, RNA, and even electrolytes in batteries.
I Introduction
Molecular dynamics simulations are a vital tool for protein structure prediction and drug designHollingsworth and Dror (2018). However, simulating the motions of large biological molecules with traditional MD methods is computationally expensive, requiring significant time and resources due to factors like system size and simulation timescaleZimmerman and Bowman (2016). Accelerating these computations is a key focus for computational chemists and biologists. Efforts are primarily directed towards two approaches: hardware and software improvements. In the hardware arena, traditional MD codes have been rewritten to leverage the parallel processing power of Graphics Processing Units (GPUs)Bergdorf et al. (2021). Additionally, the development of supercomputers like Anton has significantly boosted computational speedShaw et al. (2021). On the software front, researchers employ enhanced sampling methods, such as replica exchange methods, to explore a wider range of configurations within the simulated systemBernardi, Melo, and Schulten (2015). Additionally, coarse-grained MD simulations focus on capturing essential features of the system while averaging over less crucial details. This approach offers significant computational and conceptual advantages compared to more detailed all-atom models.
Coarse-grained methods face several challenges due to the inherent complexity of biological molecules like proteins. These challenges include capturing the intricate motions of protein backbones due to the presence of flexible side chains, loops, and disulfide bonds. Additionally, describing an implicit solvent environment that accurately reflects the crucial role of water molecules in stabilizing folded protein structures, particularly hydrophobic interactions, proves difficult with CG methods. To overcome these challenges, CG methods employ two main paradigms: "top-down" and "bottom-up".Noid (2013). The "top-down" CG method constructs collective variables (CVs) based on experimental observations, often leveraging thermodynamic data. In contrast, the bottom-up approach utilizes classical or empirical atomic models. It requires a deeper understanding of the underlying physics and chemistry. Here, researchers leverage knowledge of fundamental properties like bond lengths, bond angles, dihedral angles, and non-bonded interactions to derive a coarse-grained representation of the system.
Within the bottom-up paradigm of CG methods, researchers leverage the framework of statistical mechanics to define effective interactions, such as potential energiesLiwo et al. (2001). Over the past decades, several traditional coarse-grained methods have emerged, demonstrating significant improvements in capturing protein dynamicsMim et al. (2012); Chu and Voth (2005); Yu et al. (2021); Tóth (2007); Li and Takada (2010). The incorporation of machine learning (ML) techniques has further enhanced the performance of CG simulationsYu et al. (2021); Wang et al. (2019); Husic et al. (2020). For instance, Zhang et al. employed deep neural networks (DNNs) to approximate potential energy surfaces from coarse-grained coordinates, achieving consistent distributions of collective variables compared to those obtained from original MD simulationsZhang, Wang, and Weinan (2018). This framework further enabled enhanced sampling of large proteins Wang et al. (2022).
The three-dimensional structure of a protein, crucial for its function, is determined by a interplay between fixed covalent bonds and flexible internal angles. Among these internal angles, torsion angles (also known as dihedral angles) exhibit a more significant influence on protein conformation compared to the relatively static bond angles. Torsion angles display substantial fluctuations during MD simulations, significantly impacting how a protein folds. Consequently, these torsion angles are frequently incorporated into CG representations. Electron orbital hybridization dictates the preferred geometries of bond angles (e.g., or ). However, experimental observations and MD simulations reveal slight deviations from these ideal values. These deviations can be particularly significant when attempting to reconstruct a protein structure solely based on torsion angles. The error accumulates with each amino acid residue, as the position of an amino acid is dictated by the conformation of the preceding residues in the polypeptide chain. Therefore, CG representations that solely employ torsion angles are useful for obtaining a general understanding of a protein’s backbone conformation or predicting side-chain structures (as demonstrated in referencesZhang, Wang, and Weinan (2018); Xu, Wang, and Ma (2022); Xu et al. (2023). However, they are inadequate for high-precision tasks such as complete protein structure prediction. For instance, the AlphaFold 2 (AF2) framework utilizes torsion angles for side-chain prediction but employs Cartesian coordinates for backbone predictionJumper et al. (2021). Recent work by Kohler et al. highlights the potential benefits of incorporating bond angles. Their CG model utilizes both torsion and bond angles for the central carbon (CA) atoms, successfully capturing protein folding/unfolding transitions in small systemsKöhler et al. (2023). However, this approach does not capture the detailed dynamics of all the atoms and side chains within a protein.
In this research, we present a unified and recursive framework for constructing a bidirectional map between protein Cartesian coordinates and a minimal set of interpretable coarse-grained collective variables. This framework incorporates all heavy atoms in the backbone and side chains, enabling a comprehensive representation of the protein structure. Notably, our approach allows MD simulations to be performed solely using the collective variables, eliminating the need for explicit Cartesian coordinates. This significantly accelerates simulations while maintaining consistent performance with all-atom MD simulations, as measured by collective variable distributions and RMSDs at most time steps. Compared to existing methods that achieve only overall statistical consistency on collective variablesZhang, Wang, and Weinan (2018), our framework offers the advantage extra high accuracy in time series of RMSDs.
II Methods
II.1 Collective variables
II.1.1 coordinate transformation matrix
Our protein representation begins with a general description of atomic coordinate hierarchy in chemical compounds. For clarity, this work adopts a non-standard notation. The symbol will represent any dihedral angle and will represent any bond angles. According to referencesParsons et al. (2005), the operating matrix for rotating along a normalized rotation vector is
where is the identity matrix, and . Considering the transformation, the general formula of coordinate transformation goes as a matrix, where is a vector filled with zeros and is a matrix representing the translation.
Converting the coordinates in local axis frame (x’y’z’) to parent axis (xyz) can be understood as transforming the parent axis data (axis vector and the origin) to the the local axis data in the parent axis frame, as illustrated in Fig. 1.
This process includes three steps with operator written as
, where first, translates the frame along x axis by bond length where and ; second, rotates the frame along the new z axis by bond angle with , and ; third, rotates the frame along the new x axis by dihedral angle with , and .
For complex molecules like proteins, multiple coordinate frames are used. One frame serves as the "global" reference, and data from other frames are transformed to this reference frame. The operator for obtaining the coordinates in current local frame to the given frame is denoted as
with index showing the coordinate frames before current frame, and and being the target and current frame index. We denote the global operator for converting current coordinates to the "global" reference as
This approach leverages a recursive formula to efficiently compute the global transformation matrix for each reference frame. This eliminates the need for repeated calculations on individual sub-coordinates, simplifying the computational process. Consequently, equation
can be readily applied to determine any protein atom’s Cartesian coordinates within the global reference frame, where vector and are original Cartesian coordinates () and extended coordinates (). For convenience, subscript indicates the number and index of coordinate transformation from the "global" coordinate. A superscript of "0" denotes a "local" coordinate within its frame, otherwise it’s a "global" coordinate. As are only determined by the amino type, the Cartesian coordinates are only dependent on . The bond lengths, , used in the calculations can also be retrieved from a the standard amino acid structures. It is noteworthy that the connections between amino acids in a protein’s backbone, specifically between the nitrogen and carbon atoms, have a fixed length of about 1.32 angstroms.
The conversion between a protein’s 3D structure (Cartesian coordinates) and its internal torsion and bond angles is straightforward and is neglected here. This establishes a bidirectional mapping, facilitating analysis from both perspectives. To optimize computational efficiency, calculations consider only the essential heavy atoms. Additionally, rigid moieties like benzene rings require only two angles for complete description. This significantly reduces the number of variables needed, streamlining computational demands. While introducing some additional parameters compared to traditional methods, our approach offers substantial advantages. It allows reconstruction of the complete 3D structure, including side chains, leading to more accurate force field predictions – crucial for molecular simulations.
II.1.2 Tree data structure
Recursive data structures naturally lend themselves to representation using tree structures. This approach has been successfully implemented in J. Zhu’s prior work on quantum physics Zhu (2021); Zhu and Scrinzi (2020); Zhu (2020). Similarly, the coordinates of each reference frame within a protein can be effectively represented as a tree structure, as demonstrated in Fig. 1. In this tree, each data node inherits its information from its parent node. This tree-based representation not only captures the hierarchical relationships between protein components but also partially incorporates the protein’s geometric information. For a more detailed breakdown of the hierarchy within each amino acid, please refer to Table 1 in the appendix.
While the recursive multiplications of matrices can be computationally expensive, the reconstruction of protein 3D structures is only required during trajectory analysis. This means these computationally intensive operations are not necessary for real-time molecular dynamics simulations, thereby minimizing their impact on overall performance. Furthermore, the reconstruction of collective variable trajectories can be efficiently parallelized across CPUs and GPUs. This distributed processing approach effectively mitigates the performance drawbacks associated with recursive computation time.
II.2 Force field and molecular dynamics
Simulating large biomolecular systems necessitates capturing their statistical behavior. There exist two main approaches exist: stochastic methods and equilibrium sampling. Stochastic methods introduce random noise at each simulation step to mimic statistical fluctuations, starting from a defined initial state. Examples include Monte Carlo, random walk, and diffusion methods. Equilibrium sampling uses an initial state that already reflects the desired statistical properties, followed by noise-free simulations. This approach is employed in traditional MD simulations. Our work aligns with the second strategy (equilibrium sampling) by utilizing a machine learning force field trained on MD outputs. In contrast to methods that learn the energy landscape from existing trajectories, our approach directly predicts the next MD coordinate. This is possible because the statistical uncertainty is already embedded within the initial velocity distribution used during training. Subsequent steps involve solving Newton’s equations of motion, efficiently leveraging the learned velocity information. The network’s ability to predict the next coordinate while capturing statistical behavior becomes evident in the loss function design (Eq. 13, ). This design trains the network on a sequence of consecutive frames, effectively capturing velocity and force field information from training data. Consequently, the task becomes one of predicting the next CVs based on the previous step, as the network has already learned the underlying statistical dynamics.
For all Cartesian coordinates transformation, Eq. 6 can be written as
for N Cartesian coordinates with their corresponding constant local coordinates , M pairs of collective variables , and the overall operator being a combination of all the global operators . The inverse transformation for obtaining all the angles is written as
To overcome the periodicity of angles, in the network we project angles to sine-cosine values as
The network’s task is now to predict the subsequent collective variables based on the current ones. These predicted variables are also projected to sine-cosine values, formulated as
, where denotes the number of hidden layers in the network, and represent the time indices (timestamps) of the trajectory data. The formula of function in each layer is written as
, where and are weights to be learnt, and are activation functions. Our network distinguishes itself from Zhang et.al’s Zhang, Wang, and Weinan (2018) work by introducing a normalization constraint,, denoted by , for the sine-cosine operation within the network. This normalization step helps mitigate potential artifacts arising from the projection of angular data into sine and cosine components. While these constraints could also be incorporated into the loss function, we found that applying them directly through significantly accelerates the convergence of the network during training.
The task is to obtain the weights that minimize the loss that can be described as . We use mean square error as the loss function which goes as
, where we use error between variables at and those obtained from . This loss can also be extended to a more general form as
, where the error computed from subsequent coordinates after are also used. It is easy to see Eq. 13 collapses to Eq. 12 with . Interestingly, when in Eq. 13, the network effectively learns the underlying physical form of the force field. The velocity-verlet algorithm, a popular method in molecular simulations, allows for efficient calculation of forces acting on particles. This method uses the particles’ positions at three time steps (t-1, t, and t+1) as detailed in
The calculation works regardless of the particles’ initial velocities. In Eq. 14, is the piece-wise multiplication operator, represents a constant mass assigned to each simplified unit (coarse-grained coordinate), is the predefined time step, and the acceleration vector is the only unknown variable to be solved for. Thus including data from at least three successive trajectories enables the network to learn the force field information.
We employed two key strategies to assess the quality of the generated MD trajectories. First, we analyzed the statistical performance of all the CVs, similar to many other studies. This involved plotting the error bars associated with these CVs. Second, we quantified the overall structural similarity between the generated and reference trajectories by calculating the RMSD difference () between the predicted and reference RMSD values at each time step (t)
The reference RMSD () was obtained from the Cartesian coordinates.
III Application
III.1 Protein Structure reconstruction
To evaluate the effectiveness of our coarse-grained methods, we selected a challenging protein target: the 168-amino acid protein T1027 (sequence: KPTENNEDFNIVAVASNFATTDLDADRGKLPGKKLPL-EVLKEMEANARKAGCTRGCLICLSHIKCTPKMKKFIP-GRCHTYEGDKESAQGGIGEAIVDIPAIPRFKDLEPME-QFIAQVDLCVDCTTGCLKGLANVQCSDLLKKWLPQR-CATFASKIQGQVDKIKGAGGD) from the CASP14 prediction setsMoult (2005); Moult et al. (1995). This protein presents a computational challenge for traditional MD simulations due to its relatively long sequence and the presence of free loops. Successfully predicting the motion of this protein using our coarse-grained approach would provide a robust validation of our methods’ ability to handle complex protein dynamics.
Our coarse-grained representation utilizes 1624 parameters (described in Section II). This includes 812 bond angles and 812 dihedral angles, both on the backbone and side-chains. Compared to the 7665 parameters needed for a single protein’s Cartesian coordinates and the significantly larger number required for full MD simulations (presented in Section III.2), our approach offers a substantial reduction in dimensionality. The total number of CVs is larger than the representation using all dihedral angles for backbones that requires 504 parameters (3 angles/residue * 168 residues). This is because we include the corresponding bond angles and related side chain information. This additional complexity allows for a more precise and bidirectional representation of the protein structure.
Fig. 3 depicts the reconstructed initial structure of protein T1027. The reconstructed tertiary structure (cyan) exhibits a near-native match with the original data (gray), highlighting the accuracy of our approach. For comparison, the blue structure was built using all angles except for the bond angles, which were fixed at their theoretical value (). This comparison reveals a significant mismatch, including the absence of part of the alpha helix. This demonstrates the substantial impact of even small variations in bond angles on protein structure, emphasizing the importance of including them in our representation. The secondary structure representation also shows good agreement between the reconstructed structure and the original data, with only minor deviations observed at the ends of a few amino acid side chains. Specifically, analysis of CA atom distances reveals a maximum deviation of 0.26 angstrom, with an average value of 0.04 angstrom. The maximum difference for side-chain atoms is 0.6 angstrom, with a mean of 0.26 angstrom. It is noteworthy that we utilize local coordinates (denoted as - see Section II.1) for each amino acid, retrieved from a separate database independent of the T1027 structure. This approach can introduce slight error accumulation along the protein chain backbone, which can then propagate to the side chains (whose positions are highly dependent on the backbone). However, as will be shown in subsequent sections, the errors introduced by our coarse-grained representation remain consistently below 0.2 angstrom in practical applications, making them effectively negligible.
III.2 MD trajectories prediction
We employed the GROMACSVan Der Spoel et al. (2005); Lindahl, Hess, and van der Spoel (2001); Berendsen, van der Spoel, and van Drunen (1995) software (version 2022.2) for the MD simulations, utilizing a timestep of 1 femtosecond (fs) and saving trajectories every 0.05 nanoseconds (ns). To enhance the efficiency of the simulations, we implemented an in-house developed enhanced sampling method, parallel continuous simulated tempering and a structure based model (PCST+SBM), within GROMACS. For a detailed description of PCST+SBM, please refer to references Zhang and Ma (2010); Zang et al. (2014, 2018); Ma et al. (2018). We start the MD simulations from the refinement target in CASP14. Two copies are simulated in PCST, whose temperature distribution is defined by and , respectively. In configuring the SBM method, we selected the starting target structure as the reference to promote initial-state fidelity during the MD simulations. This ensures the dynamics primarily focus on exploring configurations near the initial structure. Compared to the original SBM implementation, we introduce a targeted constraint strategy. We constrain only the "slow" residues, which are crucial for calculating the RMSDs. Residues within loops are left unconstrained to facilitate efficient exploration of the conformational space.
For dataset preparation, we carefully select data points from the saved trajectories. Since the prediction of coordinates is inherently time-dependent, the data are sorted chronologically. Additionally, we address the artificial discontinuity introduced by replica exchange steps within the PCST methodZang et al. (2018). These exchanges, motivated by temperature differences, are not representative of the actual physical system governed by the force field. Therefore, we re-arrange the trajectories to create a continuous data stream for training. This ensures the DNN learns the underlying dynamics without being confounded by non-physical artifacts.
For training the deep neural network, we employ a common data split strategy: 70% of the processed trajectories are used for training, 20% for validation, and the remaining 10% for testing. We configure the network architecture with three hidden layers () and utilize two weight matrices () of size and (details on weight matrix dimensions might be specific to your field).
Our DNN network exhibits remarkable efficiency. It can generate a 25ns trajectory using only the initial structure as input and applying the network recursively to predict subsequent coordinates. This process takes merely 0.59 seconds on a standard NVIDIA GeForce GTX 1660 Ti (6G) GPU. In stark contrast, traditional MD simulations on a 256-CPU supercomputer require roughly 20 hours to generate a comparable trajectory. This translates to an impressive speedup of approximately 10,000 times achieved by the DNN network. Even when considering the training time (around 2 hours), the coarse-grained DNN network offers a significant speedup factor of 10 for the entire process. It’s important to note that training can be further optimized by leveraging parallelization techniques and better GPUs.
Fig. 4 depicts the predicted performance of the MD simulation. The upper panel shows the angles (torsion angles and bond angles ). As expected, the angles (represented in the first half) exhibit a wider distribution range due to their inherent flexibility in the protein backbone. In contrast, the angles (, shown in the second half) primarily cluster around 1.23 radians () with a smaller variance ( 0.4 radians) during the simulation. This finding reinforces the earlier observation of bond angle fluctuations within MD simulations. Notably, a small population of angles around radians likely originates from the carbonyl oxygen atoms in the backbone. The comparison of the orange (black) and red (predicted) curves reveals a high degree of consistency between the predicted statistical behavior of the angles and the original values. Similar observations can be made when analyzing the sine and cosine components of these angles, as presented in Fig. 5 in the appendix.
The RMSD values calculated from the predicted CVs closely match those obtained from the original Cartesian coordinates (Fig. 4, black vs. red lines). Our observations align well with previous structural analysis, where deviations were consistently less than 0.2 angstrom and the average difference using angles () was 0.0691 angstrom with a standard error of 0.0418 angstrom. Additionally, the predicted trajectories closely resemble the original data, particularly between 15-25 ns and 0-5 ns for n=2, as shown by the RMSD profiles. This overall match is also satisfied for n=5 with better agreement during 0-5 and around 12 ns, while worse agreement around 10 and 16 ns. A convergence study (Fig. 6, Appendix) explored additional n values (1, 3). As expected, n = 1 resulted in significant deviations, indicating insufficient information is learned. Increasing n to 2 eliminated these deviations, suggesting successful information capture. Interestingly, n = 3 led to a slight performance decline, while n = 5 offered minimal improvement while potentially causing slower training and overfitting. Therefore, n = 2 is optimal for convergence. The observed convergence behavior aligns with our prior analysis (Eq. 14), where appears sufficient to capture the essential force field information and generate meaningful data for MD simulations. The close agreement between our predicted and reference RMSD values suggests a more accurate representation of protein dynamics in our coarse-grained method. This is particularly noteworthy because RMSD is highly sensitive to the choice of CVs.
IV Conclusion and outlook
This research introduces a novel method using artificial intelligence to perform rapid, coarse-grained simulations of protein molecules. Our approach establishes a two-way link between simplified protein descriptions (CVs) and their detailed 3D structures. This link, built using tree-like coefficients, allows for accurate reconstruction of protein structures. Using a DNN, we successfully generated the MD trajectories of a 168-amino acid target protein (T1027) on a standard computer. This accomplishment is remarkable because it achieves speeds 10,000 times faster than traditional MD simulations, which require powerful supercomputers. Importantly, the generated trajectories closely match the statistical behavior and structural properties observed in real MD simulations. This work paves the way for highly efficient and accurate coarse-grained protein simulations using artificial intelligence.
These future research directions hold significant promise for further refining and expanding the capabilities of the presented framework.
1.
Enhanced Sampling within the Coarse-Grained Framework. While our training data incorporates enhanced sampling methods, the network may not fully capture the temperature dependence of the CV space. Future work will explore incorporating temperature and energy information directly into the DNN training process.
2.
Data Filtering and Separation of Timescales. Traditional data filtering techniques, such as those implemented in frameworks like PyEMMAScherer et al. (2015), can potentially be adapted to separate slow and fast motions within the recorded CV trajectories. The DNN could then be used to predict the slow dynamics, while fast motions could be modeled probabilistically using Gaussian noise, leading to a Monte Carlo-like simulation, as discussed in the beginning of Sec. II.2. However, further investigation is needed to determine whether the concept of "slow" and "fast" remains applicable within the CV domain.
3.
Exploration of Alternative Training Methods. Graphical Neural Networks (GNNs) offer an intriguing possibility for incorporating the inherent hierarchical structure of proteins into the training process. GNNs may be particularly well-suited for capturing the differential contributions of backbone and side-chain interactions to the overall dynamics.
4.
Expanding Applicability to Other Systems The proposed framework has the potential to be extended beyond protein simulations. By utilizing appropriate CV representations, the method could be applied to MD simulations of Deoxyribonucleic acid (DNA), Ribonucleic acid (RNA), and even non-biochemical systems such as polymers and electrolytes in batteries.
Appendix A Amino hierarchy
| Amino | 0 | 1 | 2 | 3 | 4 | 5 | 6 | ||||
| ALA | CA | CB | |||||||||
| C | O | ||||||||||
| VAL | CA | CB | CG1,CG2 | ||||||||
| C | O | ||||||||||
| LEU | CA | CB | CG | CD1,CD2 | |||||||
| C | O | ||||||||||
| GLY | CA | C | O | ||||||||
| ILE | CA | CB | CG1 | ||||||||
| CG2 | CD | ||||||||||
| C | O | ||||||||||
| MET | CA | CB | CG | SD | CE | ||||||
| C | O | ||||||||||
| TRP | CA | CB | CG |
| C | O | ||||||||||
| PHE | CA | CB | CG |
| C | O | ||||||||||
| PRO | CA | CB | CG,CD | ||||||||
| C | O | ||||||||||
| SER | CA | CB | OG | ||||||||
| C | O | ||||||||||
| CYS | CA | CB | SG | ||||||||
| C | O | ||||||||||
| ASN | CA | CB | CG | OD1,ND2 | |||||||
| C | O | ||||||||||
| GLN | CA | CB | CG | CD | OE1,OE2 | ||||||
| C | O | ||||||||||
| THR | CA | CB | OG1,CG2 | ||||||||
| C | O | ||||||||||
| TYR | CA | CB | CG |
| C | O | ||||||||||
| ASP | CA | CB | CG |
| C | O | ||||||||||
| GLU | CA | CB | CG | CD |
| C | O | ||||||||||
| LYS | CA | CB | CG | CD | CE | NZ | |||||
| C | O | ||||||||||
| ARG | CA | CB | CG | CD | NE | CZ | NH1,NH2 | ||||
| C | O | ||||||||||
| HIS | CA | CB | CG |
| C | O |
Appendix B DNN errors
Appendix C Convergence study