### v2

# A unified framework for coarse grained molecular dynamics of proteins

Abstract
Understanding protein dynamics is crucial for elucidating their biological functions. While all-atom molecular dynamics (MD) simulations provide detailed information, coarse-grained (CG) MD simulations capture the essential collective motions of proteins at significantly lower computational cost. In this article, we present a unified framework for coarse-grained molecular dynamics simulation of proteins. Our approach utilizes a tree-structured representation of collective variables, enabling reconstruction of protein Cartesian coordinates with high fidelity. The evolution of configurations is constructed using a deep neural network trained on trajectories generated from conventional all-atom MD simulations. We demonstrate the framework’s effectiveness using the 168-amino protein target T1027 from CASP14. Statistical distributions of the collective variables and time series of root mean square deviation (RMSD) obtained from our coarse-grained simulations closely resemble those from all-atom MD simulations. This method is not only useful for studying the movements of complex proteins, but also has the potential to be adapted for simulating other biomolecules like DNA, RNA, and even electrolytes in batteries.
I Introduction
Molecular dynamics simulations are a vital tool for protein structure prediction and drug designHollingsworth and Dror (2018). However, simulating the motions of large biological molecules with traditional MD methods is computationally expensive, requiring significant time and resources due to factors like system size and simulation timescaleZimmerman and Bowman (2016). Accelerating these computations is a key focus for computational chemists and biologists. Efforts are primarily directed towards two approaches: hardware and software improvements. In the hardware arena, traditional MD codes have been rewritten to leverage the parallel processing power of Graphics Processing Units (GPUs)Bergdorf et al. (2021). Additionally, the development of supercomputers like Anton has significantly boosted computational speedShaw et al. (2021). On the software front, researchers employ enhanced sampling methods, such as replica exchange methods, to explore a wider range of configurations within the simulated systemBernardi, Melo, and Schulten (2015). Additionally, coarse-grained MD simulations focus on capturing essential features of the system while averaging over less crucial details. This approach offers significant computational and conceptual advantages compared to more detailed all-atom models.
Coarse-grained methods face several challenges due to the inherent complexity of biological molecules like proteins. These challenges include capturing the intricate motions of protein backbones due to the presence of flexible side chains, loops, and disulfide bonds. Additionally, describing an implicit solvent environment that accurately reflects the crucial role of water molecules in stabilizing folded protein structures, particularly hydrophobic interactions, proves difficult with CG methods. To overcome these challenges, CG methods employ two main paradigms: "top-down" and "bottom-up".Noid (2013). The "top-down" CG method constructs collective variables (CVs) based on experimental observations, often leveraging thermodynamic data. In contrast, the bottom-up approach utilizes classical or empirical atomic models. It requires a deeper understanding of the underlying physics and chemistry. Here, researchers leverage knowledge of fundamental properties like bond lengths, bond angles, dihedral angles, and non-bonded interactions to derive a coarse-grained representation of the system.
Within the bottom-up paradigm of CG methods, researchers leverage the framework of statistical mechanics to define effective interactions, such as potential energiesLiwo et al. (2001). Over the past decades, several traditional coarse-grained methods have emerged, demonstrating significant improvements in capturing protein dynamicsMim et al. (2012); Chu and Voth (2005); Yu et al. (2021); Tóth (2007); Li and Takada (2010). The incorporation of machine learning (ML) techniques has further enhanced the performance of CG simulationsYu et al. (2021); Wang et al. (2019); Husic et al. (2020). For instance, Zhang et al. employed deep neural networks (DNNs) to approximate potential energy surfaces from coarse-grained coordinates, achieving consistent distributions of CVs compared to those obtained from original MD simulationsZhang, Wang, and Weinan (2018). This framework further enabled enhanced sampling of large proteins Wang et al. (2022).
The three-dimensional structure of a protein, crucial for its function, is determined by a interplay between fixed covalent bonds and flexible internal angles. Among these internal angles, torsion angles (also known as dihedral angles) exhibit a more significant influence on protein conformation compared to the relatively static bond angles. Torsion angles display substantial fluctuations during MD simulations, significantly impacting how a protein folds. Consequently, these torsion angles are frequently incorporated into CG representations. Electron orbital hybridization dictates the preferred geometries of bond angles (e.g., or ). However, experimental observations and MD simulations reveal slight deviations from these ideal values. These deviations can be particularly significant when attempting to reconstruct a protein structure solely based on torsion angles. The error accumulates with each amino acid residue, as the position of an amino acid is dictated by the conformation of the preceding residues in the polypeptide chain. Therefore, CG representations that solely employ torsion angles are useful for obtaining a general understanding of a protein’s backbone conformation or predicting side-chain structures (as demonstrated in referencesZhang, Wang, and Weinan (2018); Xu, Wang, and Ma (2022); Xu et al. (2023). However, they are inadequate for high-precision tasks such as complete protein structure prediction. For instance, the AlphaFold 2 (AF2) framework utilizes torsion angles for side-chain prediction but employs Cartesian coordinates for backbone predictionJumper et al. (2021). Recent work by Kohler et al. highlights the potential benefits of incorporating bond angles. Their CG model utilizes both torsion and bond angles for the central carbon (CA) atoms, successfully capturing protein folding/unfolding transitions in small systemsKöhler et al. (2023). However, this approach does not capture the detailed dynamics of all the atoms and side chains within a protein.
In this research, we present a unified and recursive framework for constructing a bidirectional map between protein Cartesian coordinates and a minimal set of interpretable coarse-grained CVs. This framework incorporates all heavy atoms in the backbone and side chains, enabling a comprehensive representation of the protein structure. Notably, our approach allows MD simulations to be performed solely using the CVs, eliminating the need for explicit Cartesian coordinates. This significantly accelerates simulations while maintaining consistent performance with all-atom MD simulations, as measured by CV distributions and RMSDs at most time steps. Compared to existing methods that achieve only overall statistical consistency on CVsZhang, Wang, and Weinan (2018), our framework offers the advantage extra high accuracy in time series of RMSDs.
II Methods
II.1 Collective variables
II.1.1 coordinate transformation matrix
Our protein representation begins with a general description of atomic coordinate hierarchy in chemical compounds. For clarity, this work adopts a non-standard notation. The symbol will represent any dihedral angle and will represent any bond angles. According to Parsons et al.Parsons et al. (2005), the operating matrix for rotating along a normalized rotation vector is
where and . Considering the transformation, the general formula of coordinate transformation goes as a matrix, where is a length-3 vector filled with zeros and is a length-3 vector representing the translation:
Converting the coordinates from the local axes frame (x’y’z’) to parent axes (xyz) can be understood as transforming the parent axes data (axes vector and the origin) to the the local axes data in the parent axes frame, as illustrated in Fig. 1.
This process includes three steps with operator written as
where first, translates the frame along x axis by bond length with with being the identity matrix, and ; second, rotates the frame along the new z axis by bond angle with , and ; third, rotates the frame along the new x axis by dihedral angle with , and . These three steps are also illustrated as ①, ② and ③ in the upper pannel of Fig. 1
For complex molecules like proteins, multiple coordinate frames are used. One frame serves as the "global" reference, and data from other frames are transformed to this reference frame. The operator for obtaining the coordinates in current local frame to the given frame is denoted as
with index showing the coordinate frames before current frame, and and being the parent and local frame index. is from Eq. 3 in the th coordinate frame. We denote the global operator for converting current coordinates to the "global" reference as
This approach leverages a recursive formula to efficiently compute the global transformation matrix for each reference frame. This eliminates the need for repeated calculations on individual sub-coordinates, simplifying the computational process. Consequently, equation
can be readily applied to determine any protein atom’s Cartesian coordinates within the global reference frame, where vector and are original Cartesian coordinates () and extended coordinates (). For convenience, subscript indicates the number and index of coordinate transformation from the "global" coordinate. A superscript of "0" denotes a "local" coordinate within its frame, otherwise it’s a "global" coordinate. As are only determined by the amino type, the Cartesian coordinates are only dependent on . The bond lengths, , used in the calculations can also be retrieved from the standard amino acid structures. It is noteworthy that the the protein backbone C-N bonds have a fixed length of about 1.32 angstroms.
The conversion between a protein’s 3D structure (Cartesian coordinates) and its internal torsion and bond angles is straightforward and is neglected here. This establishes a bidirectional mapping, facilitating analysis from both perspectives. Our calculation only consider the heavy atoms for the sake of computational efficiency. Additionally, rigid moieties like benzene rings require only two angles for complete description. This significantly reduces the number of variables needed, streamlining computational demands. While introducing some additional parameters compared to traditional methods, our approach offers substantial advantages. It allows reconstruction of the complete 3D structure, including side chains, leading to more accurate predictions of interactions – crucial for molecular simulations.
II.1.2 Tree data structure
Recursive data structures naturally lend themselves to representation using tree structures. This approach has been successfully implemented in J. Zhu’s prior work on quantum physics Zhu (2021); Zhu and Scrinzi (2020); Zhu (2020). Similarly, the coordinates of each reference frame within a protein can be effectively represented as a tree structure, as demonstrated in Fig. 1 using an artificial protein with first two amino acids being ILE and TRP as an example. This tree structure efficiently represents protein hierarchy and geometry. Each node serves as a local reference frame, storing the local coordinates of its constituent atoms. Nodes inherit the transformation operator, denoted as (refer to Eq. 5), from their parent nodes (). This inheritance reflects the hierarchical relationships within the protein. Furthermore, the tree captures geometric information. For instance, atoms in TRP forming rigid rings (e.g., CG CH2) reside within a single node. For a more detailed breakdown of the hierarchy within each amino acid, please refer to Table 1 in the appendix. This approach acknowledges their inherent rigidity and reduces the number of redundant parameters needed to define their relative positions, promoting both data storage efficiency and a more natural representation of the protein’s structure.
While the recursive multiplications of matrices can be computationally expensive, the reconstruction of protein 3D structures is only required during trajectory analysis. This means these computationally intensive operations are not necessary for real-time molecular dynamics simulations, thereby minimizing their impact on overall performance. Furthermore, the reconstruction of CV trajectories can be efficiently parallelized across CPUs and GPUs. This distributed processing approach effectively mitigates the performance drawbacks associated with recursive computation time.
II.2 Propagation and molecular dynamics
Simulating large biomolecular systems necessitates capturing their statistical behavior. There exist two main approaches: stochastic methodsHarrison (2009); G Bhanot (1988) and equilibrium samplingKarplus and Petsko (1990); Bussi and Parrinello (2007). Equilibrium sampling uses an initial state that already reflects the desired statistical properties, followed by noise-free simulationsKarplus and Petsko (1990). Stochastic methods introduce random noise at each simulation step to mimic statistical fluctuations, starting from a defined initial state. Examples include random walkCodling, Plank, and Benhamou (2008) and could be decribed by stochastic differential equations (SDE)Slavík (2013). The SDE is also widely applied in computer science for generation of figures Song and Ermon (2019); Song et al. (2021), and later applied molecular dynamics schemes to generate MD trajectoriesWu and Li (2023); Petersen, Roig, and Covino (2021).
This section explores the potential of leveraging strategically chosen data points with minimal correlation from MD trajectories to partially mimic SDEs for statistical analysisCodling, Plank, and Benhamou (2008). By constructing a dataset that reflects the system’s exploration of phase space, we can gain relative information about its dynamics through the lens of an SDE. However, careful selection of the time interval between saved points is crucial, as large gaps can miss important events and hinder accurate analysisSlavík (2013). This approach offers a potentially faster and more efficient way to gain statistical insights into complex systems, particularly when traditional methods become computationally expensive. Our work uses the strategy by utilizing a machine learning propagator trained on saved MD trajectories and this propagator directly predicts the coordinates of next snapshot. We will show that this propagator effectively includes the drift velocity in the SDE.
II.2.1 Propagator network
For all Cartesian coordinates transformation, Eq. 6 can be written as
for N Cartesian coordinates with their corresponding constant local coordinates , M pairs of collective variables , and the overall operator being a combination of all the global operators . The inverse transformation for obtaining all the angles is written as
To overcome the periodicity of angles, in the network we project angles to sine-cosine values as
The network’s task is now to predict the subsequent CVs based on the current ones. These predicted variables are also projected to sine-cosine values, formulated as
where denotes the number of hidden layers in the network, and represent the time indices (timestamps) of the trajectory data. The formula of function in th layer is written as
where and are weights to be learnt, and are activation functions. Our network distinguishes itself from Zhang et al.’s Zhang, Wang, and Weinan (2018) work by introducing a normalization constraint, denoted by , for the sine-cosine operation within the network. This normalization step helps mitigate potential artifacts arising from the projection of angular data into sine and cosine components. While these constraints could also be incorporated into the loss function, we found that applying them directly through significantly accelerates the convergence of the network during training.
The task is to obtain the weights that minimize the loss that can be described as . We use mean squared error (MSE) as the loss function which goes as
where we use error between variables at and those obtained from . This loss can also be extended to a more general form as
where the error computed from subsequent coordinates after are also used. It is easy to see Eq. 13 collapses to Eq. 12 with .
II.2.2 Propagator and SDE
The SDE describes how the CVs evolve over time. Equation
details this evolution, where where represents Gaussian white noise, is the standard deviation dependent on the CVs at time and step size , and denotes the true propagator, which corresponds to the drift velocity in Eq. 18. A subset of loss function in Eq. 13 for predicting the subsequent coordinate based on current time stamp for any time indexes can be written as
where indexes the training data, indicates the number of steps after the current time stamp and subscript refers to the true CVs. The detailed derivation of this loss function can be found in Sec. C.0.1 in the appendix. It can be seen when the loss in Eq. 13 reaches its minimum, the trained DNN effectively approximates the drift velocity represented in Eq. 14. Without the noise term, Eq. 14 collapses to Eq. 10.
Our work builds upon recent advances in SDE-based machine learning, where models learn to reverse a data corruption process (diffusion) Ho, Jain, and Abbeel (2020); Nichol and Dhariwal (2021). However, we propose a novel training scheme due to key differences in our setting. First, data characteristics: Unlike prior works that employ artificially generated, noise-driven trajectoriesHo, Jain, and Abbeel (2020); Nichol and Dhariwal (2021), our model deals with physical, sequential data with inherent physical constraints. Second, Process direction: We learn a forward process directly, contrasting with the reversed process learned in diffusion models. Third, data scarcity: We operate with limited real-world data, as opposed to the large, artificially generated datasets used in previous works. Computing the noises with limited trajectory data may include large numerical errors.
However, we present alternative ways for computing the noises. We make the approximation that at each time stamp of SDE equation, the system stays at equilibrium state and the noise part is formed by the velocities () of the CVs that satisfy the Boltzmann distribution: , where is the equivalent mass, is the Boltzmann constant and is the temperature. Due to the derivation in Sec. C.0.2 in the appendix, all the standard deviation vectors of in the dataset is proportional to a constant vector , which is only dependent on the topology of the protein. Note that the summation of all these standard deviation vectors is proportional to the loss function vector in Eq. 13. Thus we have
where we use individual elements of the vector here compared to summed elements and in Eq. 13. Thus we can use the standard deviation vector in the whole dataset as the noise vector, and select an appropriate coefficient based on the behavior of RMSD in the training data to do the simulation.
II.2.3 Model quality assessment
Due to the inherent uncertainty in the timestep of the SDE, where saved trajectories become uncorrelated after an unknown number of intervals, a range of interval sizes needs to be considered in practical applications. This choice of not only enhances the accuracy of the drift term over a larger timescale but also helps prevent the system from becoming trapped in local minima, when only the propagator operates on the SDE without the noise term for a larger interval . Interestingly, when in Eq. 13, the network demonstrably learns the underlying physical form of the drift force, leading to a more precise simulation, see Sec. D in the appendix. Therefore, we recommend using in practice. We will further substantiate this recommendation with numerical evidence in the application section.
We employed two key strategies to assess the quality of the generated MD trajectories. First, we analyzed the statistical performance of all the CVs, similar to many other studies. This involved plotting the error bars associated with these CVs. Second, we quantified the overall structural similarity between the generated and reference trajectories by calculating the RMSD difference () between the predicted and reference RMSD values at each time step (t)
The reference RMSD () was obtained from the Cartesian coordinates.
III Application
III.1 Protein Structure reconstruction
To evaluate the effectiveness of our coarse-grained methods, we selected a challenging protein target: the 168-amino acid protein T1027 (sequence: KPTENNEDFNIVAVASNFATTDLDADRGKLPGKKLPL-EVLKEMEANARKAGCTRGCLICLSHIKCTPKMKKFIP-GRCHTYEGDKESAQGGIGEAIVDIPAIPRFKDLEPME-QFIAQVDLCVDCTTGCLKGLANVQCSDLLKKWLPQR-CATFASKIQGQVDKIKGAGGD) from the CASP14 prediction setsMoult (2005); Moult et al. (1995). This protein presents a computational challenge for traditional MD simulations due to its relatively long sequence and the presence of free loops. Successfully predicting the motion of this protein using our coarse-grained approach would provide a robust validation of our methods’ ability to handle complex protein dynamics.
Our coarse-grained representation utilizes 1624 parameters (described in Section II). This includes 812 bond angles and 812 dihedral angles, both on the backbone and side-chains. Compared to the 7665 parameters needed for a single protein’s Cartesian coordinates and the significantly larger number required for full MD simulations (presented in Section III.2), our approach offers a substantial reduction in dimensionality. The total number of CVs is larger than the representation using all dihedral angles for backbones that requires 504 parameters (3 angles/residue * 168 residues). This is because we include the corresponding bond angles and related side chain information. This additional complexity allows for a more precise and bidirectional representation of the protein structure.
Fig. 3 depicts the reconstructed initial structure of protein T1027. The reconstructed tertiary structure (cyan) exhibits a near-native match with the original data (blue), highlighting the accuracy of our approach. For comparison, the gray structure was built using all angles except for the bond angles, which were fixed at their theoretical value (). This comparison reveals a significant mismatch, including the absence of part of the alpha helix. This demonstrates the substantial impact of even small variations in bond angles on protein structure, emphasizing the importance of including them in our representation. The secondary structure representation also shows good agreement between the reconstructed structure and the original data, with only minor deviations observed at the ends of a few amino acid side chains. Specifically, analysis of CA atom distances reveals a maximum deviation of 0.26 angstroms, with an average value of 0.04 angstroms. The maximum difference for side-chain atoms is 0.6 angstroms, with a mean of 0.26 angstroms. It is noteworthy that we utilize local coordinates (denoted as - see Section II.1) for each amino acid, retrieved from a separate database independent of the T1027 structure. This approach can introduce slight error accumulation along the protein chain backbone, which can then propagate to the side chains (whose positions are highly dependent on the backbone). However, as will be shown in subsequent sections, the errors introduced by our coarse-grained representation remain consistently below 0.2 angstroms in practical applications, making them effectively negligible.
III.2 MD trajectories prediction
We employed the GROMACSVan Der Spoel et al. (2005); Lindahl, Hess, and van der Spoel (2001); Berendsen, van der Spoel, and van Drunen (1995) software (version 2022.2) for the MD simulations, utilizing a timestep of 1 femtosecond (fs) and saving trajectories every 0.05 nanoseconds (ns). To enhance the efficiency of the simulations, we implemented an in-house developed enhanced sampling method, parallel continuous simulated tempering and a structure based model (PCST+SBM) within GROMACS. Using SBM focuses on the structural variations near a reference structure. This is because comprehensively analyzing all possible protein configurations is computationally demanding for large proteins, and this paper prioritizes a different aspect of the analysis. For a detailed description of PCST+SBM, please refer to references Zhang and Ma (2010); Zang et al. (2014, 2018); Ma et al. (2018). We start the MD simulations from the refinement target in CASP14. Two copies are simulated in PCST, whose temperature distribution is defined by and , respectively. The replica exchange is performed every 5 ns and the time step for updating temperature in Langevin equation is 0.01 ps. In configuring the SBM method, we selected the starting target structure as the reference to promote initial-state fidelity during the MD simulations. This ensures the dynamics primarily focus on exploring configurations near the initial structure. Compared to the original SBM implementation, we introduce a targeted constraint strategy. We constrain only the "slow" residues, which are crucial for calculating the RMSDs. Residues within loops are left unconstrained to facilitate efficient exploration of the conformational space.
For dataset preparation, we carefully select data points from the saved trajectories. Since the prediction of coordinates is inherently time-dependent, the data are sorted chronologically. Additionally, we address the artificial discontinuity introduced by replica exchange steps within the PCST methodZang et al. (2018). These exchanges, motivated by temperature differences, are not representative of the actual physical system governed by the force field. Therefore, we re-arrange the trajectories to create a continuous data stream for training. This ensures the DNN learns the underlying dynamics without being confounded by non-physical artifacts.
This coarse-grained MD simulation utilizes saved trajectories from a prior all-atom MD simulation. The all-atom simulation employed a timestep of 1 fs with data saved every 0.05 ns for a total simulation length of 250 ns. Since this study focuses on replicating the time-series behavior of the simulation, utilizing the full configuration space data becomes unnecessary. While SBM inherently doesn’t provide access to this complete data, obtaining it for such a long protein would also be computationally expensive. Our focus on time-series allows us to achieve a more precise result without requiring the full configuration space. For training the deep neural network, we employ a common data split strategy: 70% of the processed trajectories are used for training, 20% for validation, and the remaining 10% for testing. We configure the network architecture with three hidden layers () and utilize two weight matrices () of size and (details on weight matrix dimensions might be specific to your field).
Our DNN exhibits remarkable efficiency. It can generate a 25ns trajectory using only the initial structure as input and applying the network recursively to predict subsequent coordinates. This process takes merely 0.59 seconds on a standard NVIDIA GeForce GTX 1660 Ti (6G) GPU. In stark contrast, traditional MD simulations on a 256-CPU supercomputer require roughly 20 hours to generate a comparable trajectory. This translates to an impressive speedup of approximately 10,000 times achieved by the DNN network. Even when considering the training time (around 2 hours), the coarse-grained DNN network offers a significant speedup factor of 10 for the entire process. It’s important to note that training can be further optimized by leveraging parallelization techniques and better GPUs.
We first do simulations simply with our DNN propagator using Eq. 10 to show the accuracy of predicted MD trajectories using the test dataset. Fig. 4 depicts the predicted performance of the MD simulation. The upper panel shows the angles (torsion angles and bond angles ). As expected, the angles (represented in the first half) exhibit a wider distribution range due to their inherent flexibility in the protein backbone. In contrast, the angles (, shown in the second half) primarily cluster around 1.23 radians () with a smaller variance ( 0.4 radians) during the simulation. This finding reinforces the earlier observation of bond angle fluctuations within MD simulations. Notably, a small population of angles around radians likely originates from the carbonyl oxygen atoms in the backbone. The comparison of the orange (black) and red (predicted) curves reveals a high degree of consistency between the predicted statistical behavior of the angles and the original values. Similar observations can be made when analyzing the sine and cosine components of these angles, as presented in Fig. E.1 in the appendix.
The RMSD values calculated from the predicted CVs closely match those obtained from the original Cartesian coordinates (Fig. 4, black vs. red lines). Our observations align well with previous structural analysis, where deviations were consistently less than 0.2 angstroms and the average difference using angles () was 0.0691 angstroms with a standard error of 0.0418 angstroms. Additionally, the predicted trajectories closely resemble the original data, particularly between 15-25 ns and 0-5 ns for n=2, as shown by the RMSD profiles. This overall match is also satisfied for n=5 with better agreement during 0-5 and around 12 ns, while worse agreement around 10 and 16 ns. A convergence study (Fig. E.2, Appendix) explored additional n values (1, 3). As expected, n = 1 resulted in significant deviations, indicating insufficient information is learned. Increasing n to 2 eliminated these deviations, suggesting successful information capture. Interestingly, n = 3 led to a slight performance decline, while n = 5 offered minimal improvement while potentially causing slower training and overfitting. Therefore, n = 2 is optimal for convergence. The observed convergence behavior aligns with our prior analysis (Eq. 28), where appears sufficient to capture the essential force field information and generate meaningful data for MD simulations.
We subsequently conducted simulations that included noise terms, calculated from Eq. 16, using training data sets. An analogous propagation with noise was carried out for a duration of 250 ns using Eq. 14. The resulting RMSD trajectories are illustrated in Fig. E.3 in the appendix. These trajectories exhibit significant variance in RMSD values. This variability likely originates from two sources: (1) inaccuracies associated with approximating the noise terms, and (2) the artificial noise scaling factor employed. Despite this variance, the RMSD curves consistently reach values below 3Å, indicating proximity to the native protein configuration, a conformation not explicitly present in the training data. This suggests that the method can be effective in locating the folded state of a protein, but careful calibration of the noise term formulation is crucial.
Our CG method achieves a more precise representation of protein dynamics, as evidenced by the close agreement between our predicted and reference RMSD values, which was never achieved in previous CG methodsWang et al. (2022); Zhang, Wang, and Weinan (2018); Yu et al. (2021); Wang et al. (2019); Husic et al. (2020). The superior precision stems from two key features. First, incorporating both bond and dihedral angles allows for near-native reconstruction of the protein backbone, significantly improving its structural representation. Second, including side-chain information enhances the precision of interactions between amino acids, leading to more realistic protein behavior. Furthermore, the ability to perform bi-directional conversion between CG representation and all-atom structure empower our CG simulations to function as a truly independent tool, extending beyond a complementary role to traditional all-atom simulationsWang et al. (2022); Zhang, Wang, and Weinan (2018).
IV Conclusion and outlook
This research introduces a novel method using artificial intelligence to perform rapid, coarse-grained simulations of protein molecules. Our approach establishes a two-way link between simplified protein descriptions (CVs) and their detailed 3D structures. This link, built using tree-like coefficients, allows for accurate reconstruction of protein structures. Using a DNN, we successfully generated the MD trajectories of a 168-amino acid target protein (T1027) on a standard computer. This accomplishment is remarkable because it achieves speeds 10,000 times faster than traditional MD simulations, which require powerful supercomputers. Importantly, the generated trajectories closely match the statistical behavior and structural properties observed in real MD simulations. This work paves the way for highly efficient and accurate coarse-grained protein simulations using artificial intelligence.
These future research directions hold significant promise for further refining and expanding the capabilities of the presented framework.
1.
Enhanced Sampling within the Coarse-Grained Framework. While our training data incorporates enhanced sampling methods, the network may not fully capture the temperature dependence of the CV space. Future work will explore incorporating temperature and energy information directly into the DNN training process.
2.
Exploration of Alternative Training Methods. Graphical Neural Networks (GNNs) offer an intriguing possibility for incorporating the inherent hierarchical structure of proteins into the training process. GNNs may be particularly well-suited for capturing the differential contributions of backbone and side-chain interactions to the overall dynamics.
3.
Expanding Applicability to Other Systems The proposed framework has the potential to be extended beyond protein simulations. By utilizing appropriate CV representations, the method could be applied to MD simulations of Deoxyribonucleic acid (DNA), Ribonucleic acid (RNA), and even non-biochemical systems such as polymers and electrolytes in batteries.
Appendix A Example amino structure
Appendix B Amino hierarchy
| Amino | 0 | 1 | 2 | 3 | 4 | 5 | 6 | ||||
| ALA | CA | CB | |||||||||
| C | O | ||||||||||
| VAL | CA | CB | CG1,CG2 | ||||||||
| C | O | ||||||||||
| LEU | CA | CB | CG | CD1,CD2 | |||||||
| C | O | ||||||||||
| GLY | CA | C | O | ||||||||
| ILE | CA | CB | CG1 | ||||||||
| CG2 | CD | ||||||||||
| C | O | ||||||||||
| MET | CA | CB | CG | SD | CE | ||||||
| C | O | ||||||||||
| TRP | CA | CB |
| C | O | ||||||||||
| PHE | CA | CB | CG |
| C | O | ||||||||||
| PRO | CA | CB | CG,CD | ||||||||
| C | O | ||||||||||
| SER | CA | CB | OG | ||||||||
| C | O | ||||||||||
| CYS | CA | CB | SG | ||||||||
| C | O | ||||||||||
| ASN | CA | CB | CG | OD1,ND2 | |||||||
| C | O | ||||||||||
| GLN | CA | CB | CG | CD | OE1,OE2 | ||||||
| C | O | ||||||||||
| THR | CA | CB | OG1,CG2 | ||||||||
| C | O | ||||||||||
| TYR | CA | CB | CG |
| C | O | ||||||||||
| ASP | CA | CB | CG |
| C | O | ||||||||||
| GLU | CA | CB | CG | CD |
| C | O | ||||||||||
| LYS | CA | CB | CG | CD | CE | NZ | |||||
| C | O | ||||||||||
| ARG | CA | CB | CG | CD | NE | CZ | NH1,NH2 | ||||
| C | O | ||||||||||
| HIS | CA | CB | CG |
| C | O |
Appendix C Physical property of propagation
C.0.1 Drift velocity
In this section, we aim to demonstrate that the propagation operator serves as the driving force in the SDE. The SDE is typically expressed as
where represents the position within the system’s phase or state space at time . Here, is a flow vector field or drift force that signifies the deterministic evolution law, while is a collection of vector fields that define the system’s coupling to Gaussian white noise Slavík (2013). is a variable from a Gaussian distribution possessing a mean of and a variance of . By employing discrete time stamps with a constant interval of , the position at the subsequent time stamp is
In this case, is used to represent the true drift force part, and is the Gaussian noise. When expressed in the form of CVs, the value at the next time stamp, , is derived from the previous time stamp, :
Similarly, for the following two steps can be represented as
Further derivation demonstrates that the coordinates at any arbitrary time step, denoted as , can be formulated as the drift force derived from and a Gaussian noise with a mean value of zero:
This implies that the drift velocity of a larger time interval SDE can be viewed as a sequential operation of drift velocities over smaller time intervals.
The MSE, used for predicting the subsequent coordinate based on the current time stamp for any time indexes , can be expressed as
where symbols and represent the index of the training data and the steps following the current timestamp, respectively. It’s straightforward to observe that the learned network becomes the drift force once the loss achieves its minimum value.
C.0.2 Noise
We approximate that for every timestamp in the SDE, the system remains in a state of equilibrium. Here, the noise component is constituted by the velocities () of the control variables, which adhere to the Boltzmann distribution. The probability distribution for the velocities is given by . In this equation, symbolizes the equivalent mass, stands for the Boltzmann constant, and represents the temperature. The standard deviation vector from the noise vector in Eq. 22 can be expressed as
is dependent on time, while the vector , representing the equivalent mass, only depends on the structure’s topology. Each element within this vector signifies the mass, which remains unchanged throughout the simulation. From Eq. 23, the residues of a specified data sets satisfy
This suggests that the standard deviation of all existing time points adheres to
It’s straightforward to derive
which is the vector included in the loss function given by Eq. 13. Please note that this reveals we can utilize the existing training data to compute the residues. Each dimension of the CV’s residues, with a suitable scaling factor, can act as the noise of SDE in the future propagation process. This scaling factor, however, is unknown and constantly changes over time, a result of the PCST strategy applied to the training data. Despite this, we can still employ a trial-and-error strategy to choose a suitable and constant parameter (constant temperature simulation) based on the RMSD performance in relation to the original data. With this parameter, we are able to conduct a simulation in conjunction with the learned drift velocity network and the standard deviation vector.
Appendix D Force-field information
We show here, when in Eq. 13, the network effectively learns the underlying physical form of the force of the drift, which gives a more precise simulation. We illustrate this by the velocity-verlet algorithm, which is a popular method in molecular simulations and allows for efficient calculation of forces acting on particles. This method uses the particles’ positions at three time steps (t-1, t, and t+1) as detailed in
The calculation works regardless of the particles’ initial velocities. In Eq. 28, is the piece-wise multiplication operator, represents a constant mass assigned to each simplified unit (coarse-grained coordinate), is the predefined time step, and the acceleration vector is the only unknown variable to be solved for. We would like to note that Eq. 28 is only an illustration to show the inclusion of force-field information with where many details are excluded. For example, the much larger time step than the MD simulation indicates the computation of force is not accurate in the equation and the dependence of the velocity of the time due to our PCST method indicates the force should be dependent of time. However, this equation helps indicate that least three successive trajectories enables the network to learn the drift force information.
Appendix E Propagation simulation